OpenStudy (anonymous):

find the volume of the solid bounded by z=-y, the xy plane, and r=4cos(theta)

OpenStudy (anonymous):

\[\int\limits_{-\pi/2}^{0}\int\limits_{0}^{4\cos(\theta)}\int\limits_{0}^{-\sin(\theta)}r dz dr d(\theta)\]

OpenStudy (anonymous):

my question is: 4cos(theta) is a circle of radius 2 , when i integrate with that in approach \[\int\limits_{-\pi/2}^{0}\int\limits_{0}^{2}\int\limits_{0}^{-\sin(\theta)}r dz dr d(\theta)\]

OpenStudy (anonymous):

i am off by a factor of 2, am i doing something wrong!!!

OpenStudy (anonymous):

hmmm

OpenStudy (anonymous):

Can you draw it?

OpenStudy (anonymous):

ill try

OpenStudy (anonymous):

No, they are z, r, theta

OpenStudy (anonymous):

It's called cylindrical coordinates.

OpenStudy (anonymous):

yeah i tried.. no dice

OpenStudy (anonymous):

@chris2k59 Did you try to integrate normally?

OpenStudy (anonymous):

what do you mean ?

OpenStudy (anonymous):

I can't see how integrating over this will work.\[ \int\limits_{-\pi/2}^{0}\int\limits_{0}^{2}\int\limits_{0}^{-\sin(\theta)}r dz dr d(\theta) \]

OpenStudy (anonymous):

So why not integrate over the actual setup?

OpenStudy (anonymous):

r=4cos(theta) r^2=4rcos(theta) r^2=4x x^2+y^2=4x then by completing the square (x-2)^2+y^2=4; r=2

OpenStudy (anonymous):

the original setup works fine, but when used this approach i was off by a factor of 2... i want to know if i just had a mistake somewhere

OpenStudy (anonymous):

is this approach valid???

OpenStudy (anonymous):

Unfortunately your approach is not valid because an offset circle is going to have a different domain than a centered circle.

OpenStudy (anonymous):

You are changing the domain, which in general fails.

OpenStudy (anonymous):

lets make the region first start with "z"

OpenStudy (anonymous):